### v1

@@ 1 Introduction
is fundamental in statistics, scientific computation, and machine learning. Its solution is conveniently expressed using the Moore–Penrose pseudoinverse, $\boldsymbol{x}_{\star}=\boldsymbol{A}^{+}\boldsymbol{b}$. However, computing this solution is slow and memory-intensive when the number of rows is large. For the largest problems (say, $n\geq 10^{12}$), storing even a single column of $\boldsymbol{A}$ in random-access memory is challenging.
@@ 1.1 Randomized Kaczmarz
RK can be interpreted as an optimized version of stochastic gradient descent for linear least-squares problems [20] that uses nonuniform selection probabilities to improve the convergence rate and eliminate the need for step size tuning. These probabilities can be precomputed using a single pass through the matrix $\boldsymbol{A}$. However, sometimes this initial computation can be avoided by using rejection sampling [20, Sec. 3]. Alternatively, RK can be implemented with uniform sampling, which is equivalent to applying RK to the diagonally reweighted least-squares problem $\min_{\boldsymbol{x}}\left\|\boldsymbol{D}\boldsymbol{b}-(\boldsymbol{D}\boldsymbol{A})\boldsymbol{x}\right\|^{2}$ for $\boldsymbol{D}=\operatorname{diag}(1/\left\|\boldsymbol{a}_{i}\right\|)$.
@@
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration Eq. 2 converges at an exponential rate, up to a finite horizon related to the inconsistency:
@@
  • Randomized extended Kaczmarz [28] accesses and manipulates the columns of $\boldsymbol{A}$ to achieve exponential convergence to $\boldsymbol{x}_{\star}$, even in the inconsistent case. However, the column manipulations are prohibitively expensive for the largest problems.
@@ 1.2 Tail-averaged randomized Kaczmarz
which depends on the burn-in time $t_{\rm b}$ and the final time $t$. Tail averaging is frequently applied in Markov chain Monte Carlo [17] to obtain a convergent estimator from stochastically varying samples. Tail averaging has also been combined with numerical optimization methods [3, Thm. 3.2]. In particular, tail averaging leads to the optimal $\mathcal{O}(1/t)$ convergence rate for stochastic gradient descent for strongly convex loss functions [22, 24, 14].
@@
Our main proposal is tail-averaged randomized Kaczmarz (TARK), which outputs the tail average Eq. 3 of the standard RK iterates Eq. 2; see Algorithm 1. Additionally, a variant of TARK for ridge regression problems is presented in Section 3.
TARK converges to the exact least-squares solution $\boldsymbol{x}_{\star}$, with no finite horizon, for both consistent and inconsistent least-squares problems:
@@
Similar to MCMC error bounds, Theorem 1.2 decomposes the mean square error into the sum of a bias term that decays exponentially fast in the burn-in time $t_{\rm b}$ and a variance term that decays as $1/t$ in the final time $t$. Thus, TARK eliminates the finite convergence horizon and converges for all inputs $\boldsymbol{A}$ and $\boldsymbol{b}$. The proof of Theorem 1.2 appears in Section 2.1.
@@ 2.1 Proof of main theorem
The proof of Theorem 1.2 follows the pattern of analysis initiated in [26], but it takes a step further by bounding the inner product terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{t+s}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{t}-\boldsymbol{x}_{\star})\bigr{]}$, which decay exponentially fast with $s$. For ease of reading, the analysis is presented as three lemmas followed by one main calculation.
@@
Iterating this equation yields the main result.
@@
Proof 2.7 (Proof of Theorem 1.2).
Start by decomposing the mean square error as follows:
@@
which completes the proof.
@@ 2.2 Optimal row-access methods
The fastest possible convergence rate for an algorithm that accesses $t$ elements of $\boldsymbol{b}$ is $\mathcal{O}(1/t)$ [6]. This lower bound places a limit on the convergence rate for any method that accesses row–entry pairs $(\boldsymbol{a}_{i},b_{i})$, and this limit is achieved by TARK. Yet the prefactor in TARK’s convergence rate is not optimal, since it depends on the square Demmel condition number $\kappa_{\rm dem}^{2}$. As a lower bound, Appendix A provides examples of challenging least-squares problems where no algorithm can guarantee error smaller than $d/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$. Compared to this bound, TARK achieves a slower convergence rate $2\kappa_{\rm dem}^{2}/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$.
Looking forward, there are several paths to improving the performance of row-access methods. First, preconditioning strategies can be used to reduce the prefactor $\kappa_{\rm dem}^{2}$ toward the theoretical minimum value of $d$; see Appendix B. Second, row-access methods can be amplified by using block-wise strategies to process several rows simultaneously. Several block-wise strategies have been proposed [21, 10, 18, 14], but it is unclear which strategy of this type is most efficient.
@@ 3 Ridge regression
    and TARK is also an option for solving this system. However, this approach with either RK or TARK adds noise to the output by treating the regularization term $\lambda\left\|\boldsymbol{x}\right\|^{2}$ stochastically; see Section 3.2 for further discussion.
@@
Compared to the error bounds for RK, the regularization plays a key role in speeding up the convergence and controlling the size of the horizon. The proof of Theorem 3.1 can be found in Appendix C.
@@ 3.1 Tail-averaged randomized Kaczmarz for ridge regression
The proof of Theorem 3.2 can be found in Appendix C. See Table 2 for a comparison of TARK-RR with other RK-based approaches.
Method Final rate Handling of $\lambda\left\|\boldsymbol{x}\right\|^{2}$ Length-$d$ vectors
@@ Acknowledgments
We thank Haoxuan Chen, Jamie Haddock, Jiang Hu, Lin Lin, Anna Ma, Christopher Musco, Deanna Needell, Kevin Shu, Joel Tropp, and Roman Vershynin for helpful conversations.
Appendix A Lower bounds
@@
Proposition A.1 (Lower bound on mean square error).
Fix $\varepsilon>0$ and $d\geq 1$. Any algorithm that can solve all least-squares problems involving all matrices $\boldsymbol{A}\in\mathbb{R}^{n\times d}$ and vectors $\boldsymbol{b}\in\mathbb{R}^{m}$ with mean square error
@@
Now suppose a least-squares solver only accesses a subset of the entries of $\boldsymbol{b}_{1}$ with index set $\mathsf{S}_{1}\subseteq\{1,\ldots,m\}$, a subset of the entries of $\boldsymbol{b}_{2}$ with index set $\mathsf{S}_{2}\subseteq\{1,\ldots,m\}$, and so on. Conditional on the revealed entries, the vector $\bigl{[}\begin{smallmatrix}\boldsymbol{b}_{i}\\ y_{i}\end{smallmatrix}\bigr{]}$ now has a conditional Gaussian distribution. In particular, applying the mean and variance formulas for a conditional Gaussian distribution yields $y_{i}\sim\mathcal{N}(m_{i},v_{i}^{2})$ for
@@
To simplify this mean square error formula, observe that as long as $\sigma^{2}>1/m$, the function $|\mathsf{S}_{i}|\mapsto-\sigma^{4}|\mathsf{S}_{i}|/\bigl{[}|\mathsf{S}_{i}|(\sigma^{2}-\frac{1}{m})+1\bigr{]}$ is convex. Since $\frac{1}{d}\sum_{i=1}^{d}|\mathsf{S}_{i}|\leq t/d$, where $t$ is the maximum number of entries accessed, it follows
@@
Proposition A.1 also leads to a bound on the mean square residual error $\operatorname{\mathbb{E}}\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}$.
Corollary A.3 (Lower bound on mean square residual error).
@@
Suppose there is an algorithm that guarantees Eq. 12 after accessing just $t<d/\varepsilon$ entries of $\boldsymbol{b}$. Then it follows that $\boldsymbol{A}^{+}\boldsymbol{A}\widehat{\boldsymbol{x}}$ is a highly accurate least-squares estimator, which satisfies
@@
But this contradicts the results of Proposition A.1, so there cannot be such an algorithm.
Appendix B Achieving the lower bounds: Preconditioning and initialization
Comparing the TARK error bound Theorem 1.2 to the lower bound Proposition A.1, we recognize two possible areas for improvement: the presence of the square Demmel condition number $\kappa_{\rm dem}^{2}$ in place of the dimension $d$ and the burn-in period needed to wash out the influence of the initialization $\boldsymbol{x}_{0}$. The former problem can by addressed by applying TARK to a preconditioned version of the least-squares problem

### v2

@@ 1 Introduction
is fundamental in statistics, scientific computation, and machine learning. Its solution is conveniently expressed using the Moore–Penrose pseudoinverse, $\boldsymbol{x}_{\star}=\boldsymbol{A}^{+}\boldsymbol{b}$. However, computing this solution by direct means is slow and memory-intensive when the number of rows is large. For the largest problems (say, $n\geq 10^{12}$), storing even a single column of $\boldsymbol{A}$ in random-access memory is challenging.
@@ 1.1 Randomized Kaczmarz
RK can be interpreted as an optimized version of stochastic gradient descent for linear least-squares problems that uses nonuniform selection probabilities to improve the convergence rate and eliminate the need for step size tuning [20]. These probabilities can be precomputed using a single pass through the matrix $\boldsymbol{A}$, which might be expensive. Sometimes this initial computation can be avoided by using rejection sampling [20, Sec. 3]. Alternatively, RK can be implemented with uniform sampling, which is equivalent to applying RK to the diagonally reweighted least-squares problem $\min_{\boldsymbol{x}}\left\|\boldsymbol{D}\boldsymbol{b}-(\boldsymbol{D}\boldsymbol{A})\boldsymbol{x}\right\|^{2}$ for $\boldsymbol{D}=\operatorname{diag}(1/\left\|\boldsymbol{a}_{i}\right\|)$.
@@
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration Eq. 2 converges exponentially fast until reaching a finite horizon related to the inconsistency:
@@
  • Randomized extended Kaczmarz [28] accesses and manipulates the columns of $\boldsymbol{A}$ to achieve exponential convergence to $\boldsymbol{x}_{\star}$, even in the inconsistent case. Yet, the column manipulations are prohibitively expensive for the largest problems.
@@ 1.2 Tail-averaged randomized Kaczmarz
which depends on the burn-in time $t_{\rm b}$ and the final time $t$. Tail averaging is frequently applied in Markov chain Monte Carlo [17] to obtain a convergent estimator from stochastically varying samples. Tail averaging has also been combined with numerical optimization methods [3, Thm. 3.2], and it leads to the optimal $\mathcal{O}(1/t)$ convergence rate for stochastic gradient descent for strongly convex loss functions [22, 24, 14].
@@
Our main proposal is tail-averaged randomized Kaczmarz (TARK), which outputs the tail average Eq. 3 of the standard RK iterates Eq. 2; see Algorithm 1. A variant of TARK for ridge regression problems will be presented in Section 3.
TARK converges to the exact least-squares solution $\boldsymbol{x}_{\star}$ with no finite horizon, for both consistent and inconsistent least-squares problems:
@@
The proof of Theorem 1.2 appears in Section 2.1.
Similar to MCMC error bounds, Theorem 1.2 decomposes the mean square error into the sum of a bias term that decays exponentially in the burn-in time $t_{\rm b}$ and a variance term that decays as $1/t$ in the final time $t$. In particular, TARK converges when both $t_{\rm b}$ and $t-t_{\rm b}$ go to infinity. To control both terms in this error bound, we recommend selecting $t_{\rm b}\in[t/4,t/2]$; see Appendix A for a storage-efficient implementation that ensures this condition when the final time $t$ is not known in advance.
@@
A similar proof guarantees that TARK converges when $t$ goes to infinity with $t_{\rm b}$ fixed. We have the following alternative version of Theorem 1.2:
Theorem 1.3 (Alternative TARK error bound).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. The TARK estimator satisfies the alternative error bound:
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}\leq\frac{2\kappa_{\rm dem}^{2}-1}{t-t_{\rm b}}\cdot\biggl{[}\frac{\left(1-\kappa_{\rm dem}^{-2}\right)^{t_{\rm b}}}{\kappa_{\rm dem}^{2}(t-t_{\rm b})}\cdot\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\right\|^{2}+\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}\biggr{]}.$
The proof of Theorem 1.3 appears in Section 2.1.
@@ 2.1 Proof of main theorem
The proof of Theorem 1.2 follows the pattern of analysis initiated in [26], but it takes a step further by bounding the inner product terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{t+s}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{t}-\boldsymbol{x}_{\star})\bigr{]}$ which decay exponentially fast with $s$. For ease of reading, the analysis is presented as three lemmas followed by one main calculation.
@@
Iterating this equation completes the proof.
@@
Proof 2.7 (Proof of Theorem 1.2 and Theorem 1.3).
First decompose the mean square error as follows:
@@
which completes the proof of Theorem 1.2. Theorem 1.3 is proved in a similar way, by explicitly averaging over term A also.
@@ 2.2 Optimal row-access methods
In one class of challenging least-squares problems, random noise is injected into $\boldsymbol{b}$, making it possible to establish rigorous lower bounds on how quickly any method which accesses a limited number of row–entry pairs $(\boldsymbol{a}_{i},b_{i})$ can converge in its estimates of $\boldsymbol{x_{\star}}$. In this setting, an optimal method that accesses just $t$ elements of $\boldsymbol{b}$ produces a mean square error that scales as $\mathcal{O}(1/t)$ as $t\rightarrow\infty$ [6]. More precisely, Appendix B provides a construction of challenging least-squares problems where any row-access algorithm leads to a mean square error of $d/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ or higher. Compared to this lower bound, Theorem 1.3 guarantees that TARK’s mean square error vanishes at a rate of $(2\kappa_{\rm dem}^{2}-1)/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ as $t\rightarrow\infty$. Therefore, TARK achieves the optimal $\mathcal{O}(1/t)$ scaling, but the prefactor in TARK’s mean square error bound is not optimal, since it depends on the square Demmel condition number $\kappa_{\rm dem}^{2}$.
Looking forward, there are several paths toward improving the mean square error of TARK and other row-access methods. First, preconditioning strategies can be used to reduce the prefactor $\kappa_{\rm dem}^{2}$ toward the theoretical minimum value of $d$; see Appendix C for an analysis of optimal preconditioning. Second, row-access methods can be amplified by using block-wise strategies to process several rows simultaneously. Several block-wise strategies have been proposed [21, 10, 18, 14], but it is unclear which strategy of this type is most efficient.
@@ 3 Ridge regression
    and TARK is also an option for solving this system. However, this approach with either RK or TARK leads to limited accuracy because it treats the regularization term $\lambda\left\|\boldsymbol{x}\right\|^{2}$ stochastically; see Section 3.2 for further discussion.
@@
Compared to the error bounds for RK, the regularization plays a key role in speeding up the convergence and controlling the size of the horizon. The proof of Theorem 3.1 can be found in Appendix D.
@@ 3.1 Tail-averaged randomized Kaczmarz for ridge regression
The proof of Theorem 3.2 can be found in Appendix D. See Table 2 for a comparison of TARK-RR with other RK-based approaches.
Method Final rate Handling of $\lambda\left\|\boldsymbol{x}\right\|^{2}$ Length-$d$ vectors?
@@ 3.2 Numerical demonstration
4 Conclusion
Randomized Kaczmarz has long served as a simple, explicitly analyzable algorithm that responds to the key scaling challenges of overdetermined linear least-squares problems. In addition, the detailed analysis of randomized Kaczmarz has highlighted broader opportunities to understand and improve stochastic gradient descent methods [20]. Building on this past research, the current work highlights the opportunity to incorporate tail averaging within randomized Kaczmarz to improve theoretical error bounds and practical efficiency. These results are highly encouraging regarding the use of tail averaging beyond linear least-squares. As additional contributions, this paper points toward preconditioning and block-wise arithmetic as opportunities to further speed up the performance of large-scale linear least-squares solvers, and it suggests alternating minimization as an effective technique for incorporating ridge regularization into stochastic iterative methods.
@@ Acknowledgments
We thank Haoxuan Chen, Zhiyan Ding, Michał Dereziński, Jamie Haddock, Jiang Hu, Lin Lin, Anna Ma, Christopher Musco, Deanna Needell, Kevin Shu, Joel Tropp, Roman Vershynin, and Jonathan Weare for helpful conversations.
Appendix A TARK with increasing burn-in time
To control both bias and variance, we recommend implementing TARK with a burn-in time that comprises a quarter to half of the final time, $t_{\rm b}\in[t/4,t/2]$. In practice, we may not know the final time $t$ in advance, opting to run the algorithm for as many iterations as needed for the solution to meet some error tolerance. To implement TARK in a storage-efficient manner in this setting, one can use the following storage-efficient TARK implementation that implements an increasing burn-in time $t_{\rm b}=2^{\lfloor\log_{2}(t)\rfloor-1}$. The approach is based on storing just two extra vectors $\boldsymbol{\widetilde{x}}_{\rm old},\boldsymbol{\widetilde{x}}_{\rm new}\in{}^{d}$, where $\boldsymbol{\widetilde{x}}_{\rm old}$ sums the TARK iterates from time $2^{\lfloor\log_{2}(t)\rfloor-1}$ to $2^{\lfloor\log_{2}(t)\rfloor}$ and $\boldsymbol{\widetilde{x}}_{\rm new}$ sums the TARK iterates from time $2^{\lfloor\log_{2}(t)\rfloor}$ to $t$. When the iteration time $t$ hits a power of $2$, the two vectors are updated via $\boldsymbol{\widetilde{x}}_{\rm old}\leftarrow\boldsymbol{\widetilde{x}}_{\rm new}$ and $\boldsymbol{\widetilde{x}}_{\rm new}\leftarrow\boldsymbol{0}$. At any time $t$, the TARK result can be quickly calculated using $\boldsymbol{\widetilde{x}}_{\rm old}$ and $\boldsymbol{\widetilde{x}}_{\rm new}$ as follows:
$\boldsymbol{\boldsymbol{\widetilde{x}}}_{t}=\frac{1}{t-t_{\rm b}}\sum_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s}=\frac{\boldsymbol{\widetilde{x}}_{\rm old}+\boldsymbol{\widetilde{x}}_{\rm new}}{t-t_{\rm b}}.$
See Algorithm 3 for pseudocode.
Algorithm 3 TARK with increasing burn-in time
1:Matrix $\boldsymbol{A}\in{}^{n\times d}$, vector $\boldsymbol{b}\in{}^{n}$, initial estimate $\boldsymbol{x}_{0}\in{}^{d}$, and final time $t$
2:$\boldsymbol{\widetilde{x}}_{\rm old}=\boldsymbol{0}$, $\boldsymbol{\widetilde{x}}_{\rm new}=\boldsymbol{0}$
3:for $s$ in $20,\ldots,t-$ 2 do
4:     Sample $i\sim\left\|\boldsymbol{a}_{i}\right\|^{2}/\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$
5:     $\boldsymbol{x}_{s+1}=\boldsymbol{x}_{s}+\frac{b_{i}-\boldsymbol{a}_{i}^{\top}\boldsymbol{x}_{s}}{\left\|\boldsymbol{a}_{i}\right\|^{2}}\boldsymbol{a}_{i}$
6:     $\boldsymbol{\widetilde{x}}_{\rm new}=\boldsymbol{\widetilde{x}}_{\rm new}+\boldsymbol{x}_{s+1}$
7:     if $s+1$ is a power of $2$ then
8:         $\boldsymbol{\widetilde{x}}_{\rm old}=\boldsymbol{\widetilde{x}}_{\rm new}$, $\boldsymbol{\widetilde{x}}_{\rm new}=\boldsymbol{0}$
9:     end if
10:end for
11:$t_{\rm b}=2^{\lfloor\log_{2}(t)\rfloor-1}$
12:$\boldsymbol{\overline{x}}_{t}=\frac{1}{t-t_{\rm b}}(\boldsymbol{\widetilde{x}}_{\rm old}+\boldsymbol{\widetilde{x}}_{\rm new})$
13:return $\boldsymbol{\overline{x}}_{t}$
Appendix B Lower bounds
@@
Proposition B.1 (Lower bound on mean square error).
Fix $\varepsilon>0$ and $d\geq 1$. Any algorithm that can solve all least-squares problems involving all matrices $\boldsymbol{A}\in\mathbb{R}^{n\times d}$ and vectors $\boldsymbol{b}\in\mathbb{R}^{n}$ with mean square error
@@
Now suppose a least-squares solver only accesses a subset of the entries of $\boldsymbol{b}_{1}$ with index set $\mathsf{S}_{1}\subseteq\{1,\ldots,m\}$, a subset of the entries of $\boldsymbol{b}_{2}$ with index set $\mathsf{S}_{2}\subseteq\{1,\ldots,m\}$, and so on. Conditional on the revealed entries, the vector $\bigl{[}\begin{smallmatrix}\boldsymbol{b}_{i}\\ y_{i}\end{smallmatrix}\bigr{]}$ now has a Gaussian distribution. In particular, applying the mean and variance formulas for Gaussian conditioning yields $y_{i}\sim\mathcal{N}(m_{i},v_{i}^{2})$ for
@@
To simplify this mean square error formula, observe that as long as $\sigma^{2}>1/m$, the function $|\mathsf{S}_{i}|\mapsto-\sigma^{4}|\mathsf{S}_{i}|/\bigl{[}|\mathsf{S}_{i}|(\sigma^{2}-\frac{1}{m})+1\bigr{]}$ is decreasing and convex. Since $\frac{1}{d}\sum_{i=1}^{d}|\mathsf{S}_{i}|\leq t/d$, where $t$ is the maximum number of entries accessed, it follows
@@
Proposition B.1 also leads to a bound on the mean square residual error $\operatorname{\mathbb{E}}\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}$.
Corollary B.3 (Lower bound on mean square residual error).
@@
Any algorithm that guarantees Eq. 12 also guarantees
@@
By Proposition B.1, the algorithm must allow access to $t\geq d/\varepsilon$ entries of $\boldsymbol{b}$.
Appendix C Achieving the lower bounds: Preconditioning and initialization
Comparing the TARK error bound Theorem 1.2 to the lower bound Proposition B.1, we recognize two possible areas for improvement: the presence of the square Demmel condition number $\kappa_{\rm dem}^{2}$ in place of the dimension $d$ and the burn-in period needed to wash out the influence of the initialization $\boldsymbol{x}_{0}$. The former problem can by addressed by applying TARK to a preconditioned version of the least-squares problem